OpenStudy (anonymous):

How to simplify radicals with coefficient? 3√ ̅512

OpenStudy (thefurball):

Remember the property that \[\sqrt{a} * \sqrt{b} = \sqrt{ab}\] Knowing that, you can separate 512 into two products, one that is a perfect square. So, 512 is the same as 256 * 2, and the sqrt(256) is 16. Can you figure it out from there?

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

is that \(\huge \sqrt[3]{512}\) ?

OpenStudy (anonymous):

Yes sir

OpenStudy (thefurball):

oh i misread that then

OpenStudy (anonymous):

Haha it's fine

OpenStudy (bibby):

well 512 = 64*8 and 64 = 8 * 8 and \(\sqrt[3]{a*b}=\sqrt[3]{a}*\sqrt[3]{b}\\ soooo\\ \sqrt[3]{512}=\sqrt[3]{64*8}\)

OpenStudy (thefurball):

Same concept, you just need to find factors of 512 that can be third rooted